OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84.

OpenStudy (lgbasallote):

y varies directly with x so y = kx y = -7 when x = -1 -7 = k(-1) -7 = -k k = 7 find x when y = -84 y = kx -84 = 7x x = \(-\frac{84}{7}\)